Sampler, Inc., which sells books to college bookstores and individuals, uses activity-based costing and activity-based management. The following information is available for the company's three cost pools:

Activity

Cost Driver

Cost-Driver Quantity

Percent of Cost-Driver Activity for Bookstore transactions

Percent of Cost-Driver Activity for Transactions to Individuals

Incoming Receipts

Number of Purchase orders

3,000

20%

80%

Warehousing

Number of inventory moves

8,000

60

40

Outgoing shipments

Number of shipments

18,000

30

70


Bookstore sales totaled $8,400,000, and sales to individuals amounted to $2,400,000. Costs for the three activities were: Incoming receipts, $450,000; warehousing, $520,000; and outgoing shipments, $630,000. A review of the company's activities found various inefficiencies with respect to the warehousing of textbooks (acquired for eventual sale to bookstores) and outgoing shipments to individuals. These inefficiencies resulted in an extra 500 moves and 400 shipments, respectively.

Required:
A. What is a non-value-added activity?
B. How much did non-value-added activities cost Sampler this past year?
C. Which of the two markets—sales to bookstores or sales to individuals—resulted in lower overall costs for incoming receipts, warehousing, and outgoing shipments? Evaluate these costs in both absolute dollars and as a percentage of sales. In addition, present a possible explanation for your results. Note: Exclude costs that arose from inefficient operations.
Answer #1

Part A

Non-value-added activities can be defined as activities that are either that are not necessary and dispensable or necessary but not efficient and improvable. In simple way, non-value-added activities are those which will not harm overall quality, performance, or perceived value even if they are eliminated.

Part B

Incoming receipts: $450,000 / 3,000 purchase orders = $150 per purchase order

Warehousing: $520,000 / 8,000 inventory moves = $65 per move

Outgoing shipments: $630,000 / 18,000 shipments = $35 per shipment

Warehousing: 500 moves x $65 = $32,500

Outgoing shipments: 400 shipments x $35=14,000

Total= $46,500

Part C

Sales to bookstores produced lower overall costs in both absolute dollars and as a percentage of sales. The main reason for this is that sales to individuals resulted in the sale of one or two copies per shipment and order. On the other hand, bookstore sales likely produced greater revenues and efficiencies because of the large number of texts sold per transaction.

Activity

Cost-DriverQuantity

% Bookstores

% Individuals

Driver-Quantity:Bookstores

Driver-Quantity:Individuals

Incoming receipts

3000

20%

80%

600

2400

Warehousing

8000

60%

40%

4300

3200

Outgoing shipment

18000

30%

70%

5400

12200

(8,000 moves x 60%) – 500 = 4300

(18,000 shipments x 70%) – 400=12200

Bookstores

Individuals

Incoming receipts:

600 purchase orders x $150

$90,000

2400 purchase orders x $150

$360,000

Warehousing:

4300 Moves x $65

279500

3200 Moves x $65

208000

Outgoing shipments:

5400 shipments x $35

189000

12200 shipments x $35

427000

Total Cost

$558,500

$995,000

Cost as a percentage of sales:

$558,500 $8,400,000

6.65%

$995,000 $2,400,000

41.46%